Calcium carbonate decomposes to form calcium oxide and carbon dioxide on strong heating. What is the volume of carbon dioxide released, measured at r.t.p, when 10.5 g of calcium oxide is formed?
1
Expert's answer
2015-11-17T08:47:40-0500
CaCO3 = CaO + CO2
nCO2 = nCaCO3 = mCaCO3/MWCaCO3 = 10.5 g/100.0869 g/mol = 0.1049 mol;
VCO2 = nCO2*Vm = 0.1049 mol*24 L/mol = 2.5178 L ≈ 2.52 L